A garden table and a bench cost $600 combined. The price of the garden table is three times the price of the bench. What is the cost of the bench?
Found 2 solutions by chibisan, rfer:
Answer by chibisan(131)

let the cost of the garden table be x and the bench be y so ,
x + y = 600 -(1)
x=3y -(2)
solve simultaneously , so
sub x=3y to eq (1)
(3y) + y = 600
4y = 600
y = 150 (the bench price) .

Answer by rfer(10432)

600/4=$150
the table is 3 of the 4
the bench is 1 of the 4